OpenStudy (anonymous):

Find the reference angle for -48 degrees.

OpenStudy (anonymous):

first find the SOCA -48 + 360 = 312 now find the reference if the terminal side lies on quadrant IV the formula is 360 - θ = ? 360 - 312 = 48 the reference angle is always positive one way to solve it is the -48 is lying ing quadrant 1 every ang that lies in quadrant 1 is always the same